Proposal:
2012B-0001
Target of Oportunity:
No
Time Observed:
8.75
Time Engineering:
0.0
Time lost technical:
0.25
Time lost weather:
0.0
Time lost other:
0.0
Description:
Dark Energy Survey
Primary Investigator:
Josh Frieman
Email:
frieman@fnal.gov [1]
Institution:
Fermilab
Instrument:
CTIO 4m DECam
Organization:
NOAO
Total Program Time:
9
Observing Team
DES Observers: Simon Birre (Obs1), Brian Nord (Obs2), M (Elizabeth Buckley-Geer_ (RM)
Telops: Alberto Alvarez, Hernan Tirado
Remote Support: —-
* Observing Plan
# standard bias and dome flat sequence
# evening standard stars
# run ObsTac queue
# end of the night standard star
# end of night calibrations
* Conditions
# weather in afternoon: 5:03pm (local): Clear blue skies overhead, a few scattered clouds near horizon, winds ~8 mph, 10C, ~55% humidity
# weather after dinner: 7pm (local): clear, no clouds, wind~8mph, 10C, 55% humidity
# weather after sunset: 9pm (local): clear, no clouds, haze in the East, wind~5mph, 10C, 60% humidity
# weather during night: ?pm (local): clear, no clouds
# weather summary: clear, 10C, low wind (hovered between 5 and 10)
# Seeing conditions: 0.8 (rms 0.1) and 1.1 (rms 0.05)
# Wind conditions: low most of night; had to raise blind and louvers around 0200(local).
* Accomplishments
# Afternoon: zeros done, (ugriz) dome flats done, flats checked
# Evening: Standards
## SDSSJ 0100-0000
## SDSSJ 0000-0000
## SDSSJ 2140-0000
# Survey Exposures:
## before moon-set: z,Y
## after moon-set g,r,i
## SN-S2 in early night
# End of night: Standards (0547hr local)
## SDSS J0958 – 0010
## SDSS J0843 – 0000
## MaxVis
* Problems tagged as Synopsis_Problem
## Rasicam: off
## ~10min lost to shutter controller error: reset at Telops Console by Telops
* Notes for Data Management
# Shutter controller error caused loss of exposure 372566. Obstac repeated the same field/hex on the next exposure.